If the scale on a model is 1 inch: 10 feet and a house is 2.5 inches tall, how tall would the actual house be?
Answer by jim_thompson5910(35256)

(1 inch)/(10 feet) = (2.5 inches)/(x feet)


1/10 = 2.5/x


1*x = 10*2.5


x = 25


The actual house is 25 feet tall (about 2 to 3 stories, depending how you define a story)